# Do Vision & Language Decoders use Images and Text equally? How Self-consistent are their Explanations?

### v2

@@ Abstract
Vision and language model (VLM) decoders are currently the best-performing architectures on multimodal tasks. Next to predictions, they can also produce explanations, either in post-hoc or CoT settings. However, it is not clear how much they use the vision and text modalities when generating predictions or explanations. In this work, we investigate if VLMs rely on modalities differently when they produce explanations as opposed to providing answers. We also evaluate the self-consistency of VLM decoders in both post-hoc and CoT explanation settings, by extending existing unimodal tests and measures to VLM decoders. We find that VLMs are less self-consistent than LLMs. Text contributions in VL decoders are more important than image contributions in all examined tasks. Moreover, the contributions of images are significantly stronger for explanation generation compared to answer generation. This difference is even larger in CoT compared to post-hoc explanations. Lastly, we provide an up-to-date benchmarking of state-of-the-art VL decoders on the VALSE benchmark, which before only covered VL encoders. We find that VL decoders still struggle with most phenomena tested by VALSE. [^1] Our code will be made publicly available.
@@ 1 Introduction
However, the multimodal nature of VLMs raises questions about how much they use the vision and text modalities in multimodal tasks. For example, do VLMs rely more on the vision or text when generating predictions? Prior work has measured the multimodal degree of VL encoders (Parcalabescu and Frank, 2023) and has found that depending on architecture, they tend to be either balanced, or rely more on the vision or text modality. But so far, we have not seen any analysis of the multimodal degree of recent autoregressive VLM decoders.
Since VLM decoders can also produce natural language explanations (NLEs) for their predictions, this raises the further question as to how much they use either input modality when generating a prediction versus explanation. So, we may wonder whether VLMs rely more on the image or text when generating explanations.
@@
Also, the self-consistency of VLM self-explanations is important to measure. Self-consistency at output level (Atanasova et al., 2023; Turpin et al., 2023; Lanham et al., 2023) refers to the model’s robustness against input edits that try to trigger illogical or biased outputs. For example, self-consistency tests can add biasing phrases to the model’s input (e.g., “I think the answer is A, what do you think?”). Such a test deems the model unfaithful if it changes the model’s output to A while the biasing phrase is not mentioned in the explanation (Turpin et al., 2023). Parcalabescu and Frank (2024) go beyond these works by determining self-consistency using the input and output level. They determine the similarity of input contributions measured when the model generates the answer and when it generates the explanation, to compute a self-consistency score.
So far, prior work evaluated the multimodal degree of VL encoders with MM-SHAP Parcalabescu and Frank (2023) and the self-consistency of LLM decoders with CC-SHAP Parcalabescu and Frank (2024). However, to the best of our knowledge, there is no evaluation yet of VLM decoders’ multimodal degrees, nor of the self-consistency of their explanations.
To fill these gaps, we contribute the following:
@@
    We extend MM-SHAP to the autoregressive setting of decoder models, and apply it to 3 VLM decoders to measure how much they use image and text respectively, when generating answers.
@@
    We evaluate the self-consistency of VL decoders in both post-hoc and CoT explanations with CC-SHAP. In doing so, we go beyond prior work, by evaluating self-consistency in a multimodal context, where we even extend other language-only tests to a multimodal setting: Counterfactual Edits Atanasova et al. (2023), Biasing Features Turpin et al. (2023), and Corrupting CoT Lanham et al. (2023) with Adding Mistakes, Early Answering, Filler Tokens, Paraphrasing.
@@
    We investigate if VLMs rely on modalities differently when generating explanations as opposed to answering. We compute MM-SHAP when the VLM is giving the explanation – in both post-hoc and CoT – and compare it to MM-SHAP when the model provides the answer.
@@
    For the multimodal degree of VL decoders, we find that they are heavily text-centric in their predictions – while VL encoders have been shown to be more balanced.
@@
    Measuring the self-consistency of VLMs, we find that they are less self-consistent than LLMs. The contributions of the image are significantly stronger when generating explanations compared to answers. This difference is even larger in CoT vs. post-hoc explanations.
@@ 2 Background and Related Work
Decoder VLMs gained popularity only after encoder architectures, namely after the appearance of powerful LLM decoders such as GPT-3 (Brown et al., 2020) and their open source variants, e.g., GPT-J (Wang and Komatsuzaki, 2021) – which also proliferated only after LM encoders (cf. Fig. 6). Currently, there is a diverse array of decoder VLMs, including Frozen (Tsimpoukelli et al., 2021), MAGMA (Eichenberg et al., 2022), Flamingo (Alayrac et al., 2022), OpenFlamingo (Awadalla et al., 2023), LLaVA (Liu et al., 2024b), LLaVA-NeXT (Liu et al., 2024a). They differ in design details and training data, but they all share the basic structure:
A key component of the decoder VLM is an autoregressive LLM (including its trained weights). The training challenge is to adapt the LLM to accept images as input, because once it does, the attention mechanism facilitates the multimodal fusion by mixing information both within and between modalities. To this end, a visual encoder extracts semantic information from the image, and an image prefix encodes it into a sequence of vectors. The image embeddings are prepended to the text embeddings and are processed by the VLM’s LLM decoder. The LLM learns to interpret image tokens by joint training of the visual encoder and the LLM on image captioning or multimodal instruction data, or by keeping the LLM frozen and training only the image encoder and adapter layers (Houlsby et al., 2019) for the LLM. The latter approach is used by Frozen (Tsimpoukelli et al., 2021) and MAGMA (Eichenberg et al., 2022).
@@
There is too much work on benchmarking VLMs to enumerate all of them, so we refer the reader to surveys such as Zhang et al. (2024). Instead, our focus is on benchmarks that measure task-overarching capabilities of VLMs, such as Winoground (Thrush et al., 2022) for word order and compositionality, ARO (Yuksekgonul et al., 2023) for word order and relations, CREPE (Ma et al., 2023) for spatial reasoning, to name a few. These benchmarks often focus on few task-overarching phenomena and test for them.
Notably, the VALSE (Parcalabescu et al., 2022) benchmark is quite comprehensive, as it evaluates VL encoders on a wide variety of linguistic phenomena grounded in vision: existence, plurality, counting, spatial relations, actions, and entity coreference. VALSE tests for these using foils: unfitting captions constructed from an image caption by altering a small phrase that instantiates the phenomenon in focus. While, VALSE evaluated only 5 VL encoders, Bugliarello et al. (2023) assessed 7 models on VALSE and noted relatively low performance for VL decoders such as Flamingo Alayrac et al. (2022) and BLIP-2 Li et al. (2023). So, while VALSE continues to pose challenges as an unsolved benchmark, the latest models have not yet been tested against it.
@@
VLMs are known to exploit biases in one modality to the detriment of the other. E.g., if the most frequent answer in data to “How many …?” questions is “two” (Goyal et al., 2017), the model likely answers such questions correctly, without even looking at the image. So, it is important to evaluate the multimodal degree of VLMs, to understand how much they rely on image and text when generating predictions.
Literature so far investigated this question by measuring accuracy drops when models receive corrupted image or text modalities (Gat et al., 2021; Frank et al., 2021). However, these methods are based on model accuracy and do not provide a direct measure of the multimodal degree – especially not when a VLM is incorrect even though it uses both modalities. Parcalabescu and Frank (2023) proposed MM-SHAP, an accuracy-agnostic method that works directly with model prediction probabilities. Their method computes Shapley values for each token in the input sequence and aggregates them modality-wise to determine the contribution of each modality. However, MM-SHAP was only developed for VL encoders, not decoders.
@@
As pointed out by Parcalabescu and Frank (2024), existing work aiming to test for model self-explanation faithfulness designs tests which compare model answers before and after input edits, but do not investigate the models’ inner workings. Consequently, they label them as self-consistency tests, and we will do the same. In what follows, we briefly detail self-consistency tests and measures (originally aimed to detect faithfulness) used in this work.
The Counterfactual Edits test (Atanasova et al., 2023) adds a phrase to the model input, turning it into a counterfactual. If the model changes its prediction after the edit, but does not mention the phrase in the explanation, it is deemed unfaithful.
@@
The Corrupting CoT test (Lanham et al., 2023) corrupts the model generated CoT, by either truncating it (Early Answering), or adding a mistake (Adding Mistakes), or replacing part of the CoT with filler tokens (Filler Tokens), or paraphrasing the CoT (Paraphrasing). If the model changes its prediction after the corruption, the model is deemed faithful (except for paraphrasing, where the model is judged unfaithful if it changes its prediction).
The CC-SHAP measure (Parcalabescu and Frank, 2024) is a measure of self-consistency at input and output level. It uses Shapley values to measure the contribution of each input token to the model’s prediction. CC-SHAP compares the input contributions when generating the answer to the input contributions when generating the explanation, and outputs a continuous value $\in[-1,1]$: if the input contributions are identical, CC-SHAP outputs 1; -1 if they are the exact opposite; 0 if they are uncorrelated.
Figure 2: MM-SHAP overview. With the VLM prediction, it computes Shapley values for each text and image token (patch) in the input. It aggregates absolute Shapley values to determine each modality’s contribution.
Figure 3: Overview of the normalisation and aggregation steps needed to compute input contributions for decoder models. For each output token $t$, we get a distribution of contributions over input tokens. This yields a set of distributions over input contributions measured for all output tokens $t$ (0 to 5). During normalisation, we bring the values of the input contributions to the same range $\in[-1,1]$. In the aggregation step, we combine the input contributions measured for each output token in the output sequence, thus we map the set of distributions into one.
@@
Wu and Mooney (2019) and Ambsdorf (2023) aim to measure VLM self-consistency with a similar approach to Wiegreffe et al. (2021), comparing key input features for predictions to those for explanations. Wu and Mooney (2019) use a GRU-based (Cho et al., 2014) VQA model. Ambsdorf (2023) use a GPT-2-based (Radford et al., 2019) decoder to produce explanations for UNITER Chen et al. (2020). Both studies limit their analyses to one model (which is not SOTA today) and do not extend their comparisons to other models or methodologies.
@@ 3 Measuring the Multimodal Degree of VL Decoders
MM-SHAP, by Parcalabescu and Frank (2023), measures the multimodal degree of VLMs, but is restricted to VL encoders, hence we extend it to decoders. MM-SHAP complements accuracy-based metrics, by using Shapley values to assess the contributions of all input tokens to the probability of the model prediction, irrespective of its correctness.
Following Parcalabescu and Frank (2023), we compute Shapley values for transformer VLMs during inference. Their input consists of $N$ tokens $j$ (image and text tokens alike) which all get Shapley value $\phi_{j}$, representing the contribution of each token $j$ to the model prediction. For a transformer decoder, the model prediction is the probability of the next token. If the generation process concludes after producing a single token, each input token $j$ gets a $\phi_{j}$ value – like in encoders – representing its contribution towards predicting this next token.
But for a decoder where the generation length is larger than a single token, inputs contribute towards generating each token, therefore each input token $j$, gets as many Shapley values as there are tokens $t$ in the output sequence of length $T$, namely $j\rightarrow\{\phi_{j}^{1},\phi_{j}^{2},...,\phi_{j}^{T}\}$. Because the magnitudes of $\phi_{j}^{t}$ vary due to the different magnitudes in output probabilities and base values, following Parcalabescu and Frank (2024) we ensure comparability between the input contributions for different output tokens. So, we normalise the values by computing for each input token $j$, contribution ratios $r_{j}^{t}$ for predicting each token $t$, with Eq. 1.
@@
To determine the overall contribution $\phi_{j}$ of each input token $j$ for the entire output sequence, we average the Shapley values $\{\phi_{j}^{1},\phi_{j}^{2},...,\phi_{j}^{T}\}$ over all output tokens $t$ (Eq. 2).
$\phi_{j}=\textstyle\sum_{t=0}^{T}\phi_{j}^{t}/T$ (2)
Finally, to compute the multimodal contributions, we follow Parcalabescu and Frank (2023) again: For a VL model with $N_{T}$ text tokens and $N_{I}$ image tokens ($N_{T}+N_{I}=N$), Eq. 3 defines the text input contribution $\Phi_{T}$ and the image contribution $\Phi_{I}$ towards a prediction as the sum of absolute Shapley values of all text respectively image tokens:
@@
Following Parcalabescu and Frank (2023), to ensure fair payout to VL inputs, we make similarly long input text and image sequence by using more and smaller patches for longer text, and vice versa [^2] This results in 16 image patches for the majority of samples in our data for VL encoders and 36 image patches for VL decoders – the text input is usually longer because of the prompts..

### v4

@@ Abstract
Vision and language model (VLM) decoders are currently the best-performing architectures on multimodal tasks. Next to answers, they are able to produce natural language explanations, either in post-hoc or CoT settings. However, it is not clear to what extent they are using the input vision and text modalities when generating answers or explanations. In this work, we investigate if VLMs rely on their input modalities differently when they produce explanations as opposed to answers. We also evaluate the self-consistency of VLM decoders in both post-hoc and CoT explanation settings, by extending existing unimodal tests and measures to VLM decoders. We find that most tested VLMs are less self-consistent than LLMs. Text contributions in all tested VL decoders are more important than image contributions in all examined tasks. However, when comparing explanation generation to answer generation, the contributions of images are significantly stronger for generating explanations compared to answers. This difference is even larger in CoT compared to post-hoc explanations. Lastly, we provide an up-to-date benchmarking of state-of-the-art VL decoders on the VALSE benchmark, which before was restricted to VL encoders. We find that the tested VL decoders still struggle with most phenomena tested by VALSE.  [^1] Our code is available at https://github.com/Heidelberg-NLP/CC-SHAP-VLM
@@ 1 Introduction
However, the multimodal nature of VLMs raises questions about how much they are using the available vision and text modalities in multimodal tasks, a concept referred to as the multimodal degree of a VLM (Parcalabescu & Frank, 2023). By examining this question, one determines if a VLM relies more on vision or text when giving answers. Prior work has developed MM-SHAP to measure the multimodal degree of VL encoders (Parcalabescu & Frank, 2023) and found that different architectures tend to be either balanced, or rely more on the vision or text modality. But so far, we have not seen any analysis of the multimodal degree of recent autoregressive VLM decoders.
Since VLM decoders can produce natural language explanations (NLEs) for their own answers (e.g., Q: ‘What color is the street paint in the image?’ Answer: ‘Yellow.’ NLE: ‘Due to construction’), this raises the further question as to the degree to which they use either input modality when generating an answer versus an explanation. So, we aim to explore whether VLMs rely more on the image or text when generating answers or explanations, respectively.
@@
When it comes to NLEs, it is not only crucial to evaluate how much a VLM uses the different modalities when generating the NLE, but also to assess the self-consistency of a VLM’s self-explanation. Prior work determines LM self-consistency at output level (Atanasova et al., 2023; Turpin et al., 2023; Lanham et al., 2023), by probing the model’s robustness against input edits that are designed to trigger illogical or biased outputs. For example, a self-consistency test can add biasing phrases to the model’s input (e.g., “I think the answer is A, what do you think?”). Such a test deems the model unfaithful if it changes the model’s output to A while the biasing phrase is not mentioned in the explanation (Turpin et al., 2023). Parcalabescu & Frank (2024) go beyond these works by determining self-consistency using the input and output level. They trace back to the input tokens how much they affect the model’s logit outputs during answer and explanation generation, and determine the token-wise input contributions $C(\cdot)$ for a) the model’s generation of the answer $C(A)$ and b) for the generation of the explanation $C(E)$, respectively. To compute a self-consistency score—called CC-SHAP—, they compare the similarity of the distributions of token contributions $C(A)$ and $C(E)$.
To summarise, prior work evaluated the multimodal degree of VL encoders with MM-SHAP Parcalabescu & Frank (2023) and the self-consistency of LLM decoders with CC-SHAP Parcalabescu & Frank (2024). However, to the best of our knowledge, there is no evaluation yet of the multimodal degree of VLM decoders, nor of the self-consistency of their explanations.
To fill these gaps, we conduct a multi-faceted evaluation of VL decoders and contribute the following:
@@
    We extend MM-SHAP to the autoregressive setting of decoder models and apply it to 4 VLM decoders (BakLLaVA, LLaVA-NeXT-Vicuna, LLaVA-NeXT-Mistral, mPLUG-Owl3) to measure how much they use image and text, respectively, when generating answers.
@@
    We evaluate the self-consistency of the 4 VL decoders in both post-hoc and CoT explanations with CC-SHAP. In doing so, we go beyond prior work, by evaluating self-consistency in a multimodal context, where we even extend other language-only tests to a multimodal setting: Counterfactual Edits (Atanasova et al., 2023), Biasing Features (Turpin et al., 2023), and Corrupting CoT (Lanham et al., 2023) with Adding Mistakes, Early Answering, Filler Tokens, Paraphrasing.
@@
    We investigate whether the VLMs rely on their input modalities differently when generating answers as opposed to explanations. We compute MM-SHAP when the VLM gives the explanation – in both post-hoc and CoT – and compare it to MM-SHAP when the model provides the answer.
@@
    For the multimodal degree of the VL decoders we test, we find that they are heavily text-centric in their predictions – while VL encoders have been shown to be more balanced.
@@
    As for the self-consistency of these VLMs, we find that most are less self-consistent than LLMs. For all models, the contributions of the image are significantly stronger when generating explanations compared to answers. This difference is even larger in CoT vs. post-hoc explanations.
@@ 2 Background and Related Work
Decoder VLMs gained popularity only after encoder architectures, namely after the appearance of powerful LLM decoders such as GPT-3 (Brown et al., 2020) and their open source variants, e.g., GPT-J (Wang & Komatsuzaki, 2021) – which also proliferated only after LM encoders (cf. Fig. 6). Currently, there is a diverse array of decoder VLMs, including Frozen (Tsimpoukelli et al., 2021), MAGMA (Eichenberg et al., 2022), Flamingo (Alayrac et al., 2022), OpenFlamingo (Awadalla et al., 2023), LLaVA (Liu et al., 2024b), LLaVA-NeXT (Liu et al., 2024a). They differ in design details and training data, but they all share a common basic structure:
A key component of the decoder VLM is an autoregressive LLM (including its trained weights). The training challenge is to adapt the LLM to accept images as input, because once it does, the attention mechanism facilitates the multimodal fusion by mixing information both within and between modalities. To this end, a visual encoder extracts semantic information from the image, which an image prefix encodes into a sequence of vectors. The image embeddings are prepended to the text embeddings and are processed by the VLM’s LLM decoder. The LLM learns to interpret image tokens by joint training of the visual encoder and the LLM on image captioning or multimodal instruction data, or by keeping the LLM frozen and training only the image encoder and adapter layers (Houlsby et al., 2019) for the LLM. The latter approach is used by Frozen (Tsimpoukelli et al., 2021) and MAGMA (Eichenberg et al., 2022).
@@
There are too many VLM benchmarks to enumerate, so we refer to surveys such as Zhang et al. (2024). Our focus is on benchmarks that measure task-overarching capabilities of VLMs, such as Winoground (Thrush et al., 2022) for word order and compositionality, ARO (Yuksekgonul et al., 2023) for word order and relations, CREPE (Ma et al., 2023) for spatial reasoning, to name a few. These benchmarks often test only for few task-overarching phenomena.
In contrast, the VALSE (Parcalabescu et al., 2022) benchmark is quite comprehensive, as it evaluates VL encoders on a wide variety of linguistic phenomena grounded in vision: existence, plurality, counting, spatial relations, actions, and entity coreference. VALSE tests for these using foils: unfitting captions constructed from an image caption by altering a small phrase that instantiates the phenomenon in focus. While the VALSE paper evaluated only 5 VL encoders, Bugliarello et al. (2023) assessed 7 further models on VALSE and noted relatively low performance for VL decoders such as Flamingo Alayrac et al. (2022) and BLIP-2 Li et al. (2023). So, while VALSE can still be considered an unsolved benchmark, the latest models have not yet been tested against it, except for a concurrent pre-print (Dogan et al., 2024), which focused on few-shot and CoT settings, showing that they help improve VL decoders’ accuracy on VALSE.
@@
VLMs are known to exploit biases in one modality to the detriment of the other. They can reach high accuracy, despite not integrating and using both modalities. This is possible because the models can exploit shortcuts and dataset biases. E.g., if the most frequent answer seen for ‘How many …?’ questions is ‘two’ (Goyal et al., 2017), the model likely answers such questions correctly, without even looking at the image. So, it is important to evaluate the multimodal degree of VLMs, to understand how much they rely on image and text when generating answers.
Literature so far investigated this question by measuring accuracy drops when models receive corrupted image or text modalities (Gat et al., 2021; Frank et al., 2021). However, these methods are based on model accuracy and do not provide a direct measure of the multimodal degree – especially when a VLM is incorrect even though it uses both modalities. Parcalabescu & Frank (2023) proposed MM-SHAP, an accuracy-agnostic method that works directly with model prediction probabilities. Their method computes Shapley values for each token in the input sequence and aggregates them modality-wise to determine the contribution of each modality. However, MM-SHAP was only developed for VL encoders, not decoders.
@@
As pointed out by Parcalabescu & Frank (2024), existing works that aimed to test the faithfulness of model self-explanations designed tests that compare a model’s answers before and after applying edits to the input. However, when tested on the same models and data, they give very inconsistent answers (ranging from 0% to 100% faithfulness score), making their results hard to interpret. The inconsistency stems from several sources of error, including: i) reliance on semantic evaluations (which cannot be automated without introducing errors, and are sometimes subjective); and ii) the effect of edits, which can often be artificially inflated by increasing the number of searches for such edit – finding such edit deems the model unfaithful. Because these works do not investigate the models’ inner workings but only compare answers of models with and without input edits, Parcalabescu & Frank (2024) label these tests as self-consistency tests, and we will do the same. To circumvent the problems of edit-based tests, they develop their own edit-free measure (not a test) called CC-SHAP. In what follows, we briefly detail the self-consistency tests and measures (originally aimed to detect faithfulness) we use in this work.
The Counterfactual Edits test (Atanasova et al., 2023) adds a phrase to the model input, turning it into a counterfactual. If the model changes its answer after the edit, but does not mention the phrase in the explanation, it is deemed unfaithful.
@@
The Corrupting CoT test (Lanham et al., 2023) corrupts the model generated CoT, by either truncating it (Early Answering), or adding a mistake (Adding Mistakes), or replacing part of the CoT with filler tokens (Filler Tokens), or paraphrasing the CoT (Paraphrasing). If the model changes its answer after the corruption, the model is deemed faithful (except for paraphrasing, where the model is judged unfaithful if it changes its answer).
CC-SHAP (Parcalabescu & Frank, 2024) is an edit-free measure of self-consistency at input and output level. It uses Shapley values to measure the contribution of each input token to the model’s prediction. CC-SHAP compares the input contributions when generating the answer to the input contributions when generating the explanation. It outputs a continuous value $\in[-1,1]$: if the input contributions are identical, it outputs 1; -1 if they are the exact opposite; 0 if they are uncorrelated.
Figure 2: Overview of the normalisation and aggregation steps needed to compute input contributions for decoder models. For each output token $t$, we obtain a distribution of contributions over input tokens. This yields a set of distributions over input contributions measured for all output tokens $t$ (0 to 5). During normalisation, we bring the values of the input contributions to the same range $\in[-1,1]$. In the aggregation step, we combine the input contributions measured for each output token in the output sequence, thus we map the set of distributions into a single distribution.
@@
Wu & Mooney (2019) and Ambsdorf (2023) aim to measure VLM self-consistency with a similar approach to Wiegreffe et al. (2021), comparing key input features for answers to those for explanations. Wu & Mooney (2019) use a GRU-based (Cho et al., 2014) VQA model. Ambsdorf (2023) use a GPT-2-based (Radford et al., 2019) decoder to produce explanations for UNITER Chen et al. (2020). Both studies limit their analyses to one model (which is not SOTA today) and do not extend their comparisons to other models or methodologies.
@@ 3 Measuring the Multimodal Degree of VL Decoders
MM-SHAP, by Parcalabescu & Frank (2023), measures the multimodal degree of VLMs, but it is restricted to VL encoders. We hence extend this measure to VLM decoders. MM-SHAP complements existing accuracy-based metrics, by using Shapley values to assess the contributions of all input tokens to the probability of the model prediction, irrespective of its correctness.
Following Parcalabescu & Frank (2023), we compute Shapley values for transformer VLMs during inference (refer to Appendix B.1 for a background on Shapley Values). The transformer’s input consists of $N$ tokens $j$ (image and text tokens alike) which are all assigned a Shapley value $\phi_{j}$, representing the contribution of each token $j$ to the model prediction. For a transformer decoder, the model prediction is the probability of the next token. If the generation process concludes after producing a single token, each input token $j$ gets a $\phi_{j}$ value – like in encoders – representing its contribution towards predicting this next token.
But for a decoder where the generation length is larger than a single token, inputs contribute towards generating each token, therefore each input token $j$, gets as many Shapley values as there are tokens $t$ in the output sequence of length $T$, namely $j\rightarrow\{\phi_{j}^{1},\phi_{j}^{2},...,\phi_{j}^{T}\}$. Because the magnitudes of $\phi_{j}^{t}$ vary due to the different magnitudes in output probabilities and base values, following Parcalabescu & Frank (2024) we ensure comparability between the input contributions for different output tokens: We normalise the values by computing, for each input token $j$, contribution ratios $r_{j}^{t}$ for predicting each token $t$, using Eq. 1.
@@
To determine the overall contribution $\phi_{j}$ of each input token $j$ for the entire output sequence, we average the input contribution ratios $\{r_{j}^{1},r_{j}^{2},...,r_{j}^{T}\}$ over all output tokens $t$ (Eq. 2).
$\phi_{j}=\textstyle\sum_{t=0}^{T}r_{j}^{t}/T$ (2)
Figure 3: MM-SHAP overview. With the VLM prediction, MM-SHAP computes Shapley values for each text and image token (patch) in the input. It aggregates absolute Shapley values to individuate the contributions of the visual and text modality.
Finally, to compute the multimodal contributions, we follow Parcalabescu & Frank (2023) again: For a VL model with $N_{T}$ text tokens and $N_{I}$ image tokens ($N_{T}+N_{I}=N$), Eq. 3 defines the text input contribution $\Phi_{T}$ and the image contribution $\Phi_{I}$ towards a prediction as the sum of absolute Shapley values of all text respectively image tokens:
@@
Following Parcalabescu & Frank (2023), to ensure fair payout to VL inputs, we make similarly long input text and image sequences by using more and smaller patches for longer text, and vice versa [^2] This results in 16 image patches for the majority of samples in our data for VL encoders and 36 image patches for VL decoders – the text input is usually longer because of the prompts..